a fanner has three fields of equal size and pays each of his workmen 60 pesos per day.He paid a total of 450 pesos to the workmen after each,working alone,plowed on the field.these workmen together,plowed the third field in 1 13/15 days.how many days did it take each man to plow a field alone?
Answer by stanbon(75887)

a fanner has three fields of equal size and pays each of his workmen 60 pesos per day.He paid a total of 450 pesos to the workmen after each,working alone,plowed on the field.these workmen together,plowed the third field in 1 13/15 days.how many days did it take each man to plow a field alone?
-----------------------------
Single Man DATA:
Time to plow the field alone = x days/job ; Rate = 1/x job/day
--------------------
Together DATA:
Time to plow the field together: 28/15 days/job ; Rate = 15/28 job/day
EQUATION:
rate + rate + rate = together rate
1/x + 1/x + 1/x = 15/28
3(1/x) = 15/28
1/x = 5/28
x = 28/5 = 5 3/5 days (Time each man would take to plow the field alone)
=================
Cheers,
Stan H.